OpenStudy (anonymous):

81+ -6x = 3 solve for x

OpenStudy (callisto):

Same as this: http://openstudy.com/study#/updates/4febd12de4b0bbec5cfb8ebf

OpenStudy (anonymous):

78=6x, x=13

OpenStudy (anonymous):

Wouldn't it be x= -13

OpenStudy (callisto):

81+ -6x = 3 81 - 6x = 3 -6x = 3 - 81 -6x = -78 x = -78/-6 = 13 Sorry reading something wrong just now.